Using double integral fund the area of region enclosed by √x+√y=√a and x+y=a


1
Expert's answer
2022-03-11T05:53:06-0500

Area bounded by the two curves = "\displaystyle\iint_{R}dydx".


The limits of the integral are obtained from the given equations.

"\sqrt{x} + \sqrt{y} = \sqrt{a} \implies\sqrt{y} = \sqrt{a}-\sqrt{x} \implies y = (\sqrt{a}-\sqrt{x})^2\\
x+y=a \implies y = a-x"


The limits for x are obtained from the two equations of the curve.

"\begin{aligned}
(\sqrt{a} - \sqrt{x})^{2} &= a - x\\
a - 2\sqrt{ax} + x &= a - x\\
2x &= 2\sqrt{ax}\\
4x^{2} &= 4ax \quad\text{(Squaring both sides)}\\ 
4x^{2} - 4ax &= 0\\
4x(x-a) &= 0
\end{aligned}"

which gives, "x = 0~ \&~ x=a".


Hence, y varies from "y = (\sqrt{a}-\sqrt{x})^2 \text{ to } y = a-x" and x varies from "x = 0 \text{ to } x = a."


Therefore,

"\begin{aligned}
\text{Area bounded by the curves } &= \displaystyle\int_{0}^{a}\int_{(\sqrt{a}-\sqrt{x})^2}^{a-x}dydx\\
&= \displaystyle\int_{0}^{a} y\Bigg|_{(\sqrt{a}-\sqrt{x})^2}^{a-x} dx\\
&= \displaystyle\int_{0}^{a} ((a-x)-(\sqrt{a}-\sqrt{x})^2)dx\\
&= \displaystyle\int_{0}^{a} (a - x - a - x +2\sqrt{ax})dx\\
&= \displaystyle\int_{0}^{a} ( - 2x + 2\sqrt{ax})dx\\ 
&= 2\displaystyle\int_{0}^{a} ( \sqrt{ax} - x)dx\\ 
&= 2 \Big(\sqrt{a} ~\frac{x^{\frac{3}{2}}}{(\frac{3}{2})} - \frac{x^2}{2}\Big)\Bigg|_0^a\\
&= 2 \Big(\frac{2}{3}\sqrt{a}\cdot a^{\frac{3}{2}} - \frac{a^2}{2}\Big)\\
&= 2 \Big(\frac{2a^2}{3} - \frac{a^2}{2}\Big)\\
&= \frac{a^2}{3}
\end{aligned}"